# All-electrically controlled spintronics in altermagnetic heterostructures

### v2

@@ Abstract
The recent development of altermagnetic materials, supporting spin splitting without net magnetization, opens new directions for spintronics that are fundamentally distinct from conventional ferromagnetic, antiferromagnetic, or spin-orbit coupling systems. Here we investigate spin-selective quantum transport in heterostructures composed of a normal metal and a two-dimensional $d$-wave altermagnet. We focus on two types of $d$-wave altermagnets, namely, weak and strong altermagnets that support close elliptic and open hyperbolic spin-resolved Fermi surfaces, respectively. Building on these distinct electronic structures, we propose all-electrically controlled spin filter and spin valve devices, where quantum resonant tunneling enables highly spin-polarized conductance tunable via gate voltage and interface transparency. In particular, we find that strong altermagnets allow gate-tunable full spin polarization that is robust against interface scattering and can be reversed by gate control. We further demonstrate that a double-gated spin valve electrically switches between parallel and antiparallel spin configurations, analogous to magnetic junctions but without the need for external magnetic fields. Our results establish both weak and strong altermagnets as promising platforms for realizing magnetic-field-free electrically tunable spintronic functionalities.
I Introduction
@@
Unconventional magnets have emerged as a new class of systems that lie beyond the traditional dichotomy of ferromagnets and antiferromagnets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. These so-called third type of magnets support spin-split Fermi surfaces resembling those of ferromagnets [12, 13], yet exhibiting globally vanishing magnetization due to a compensated magnetic ordering, akin to antiferromagnets [14, 15, 16]. Depending on the momentum-space symmetry of the exchange field and the underlying crystalline symmetries, unconventional magnets can be classified according to the angular character of the spin splitting, including $p$-, $d$-, $f$-, $g$-, and $i$-wave types [7, 8, 9, 17, 18]. Among these, $d$-, $g$-, and $i$-wave magnets are known as altermagnets (AMs), characterized by simultaneously broken time-reversal symmetry and spatial rotational symmetry, but preserving their joint symmetry [7, 8]. Furthermore, weak and strong AMs can be distinguished by the geometry of their closed and open Fermi surfaces, respectively. [7, 8, 19, 20, 21]. In contrast, $p$- and $f$-wave magnets break only rotational symmetry but preserve time-reversal symmetry originating from nonsymmorphic lattice symmetries [9, 10], typically exhibiting closed anisotropic Fermi surfaces.
@@
Unconventional magnets are promising platforms for spintronic devices [13, 14]. The anisotropic spin split Fermi surface offers mechanisms for giant magnetoresistance [22], spin-orbit-free anomalous Hall effects [23, 1, 24, 25, 26], spin-transfer torques [27, 28, 29], spin filtering effects [30], spin pumping effects [31], light-matter interactions [32, 33], light-induced spin density [34], non-Hermitian electronic responses [35, 36], strongly correlation in Mott insulators [37] and other novel phenomena [38, 39]. These mechanisms are expected in experimentally accessible altermagnetic materials such as RuO2 [40, 23, 25], MnTe [41, 24], Mn5Si3 [26, 42], CrSb [43], and Mn2Au [44], exhibiting novel spin transport. In addition to their normal-state properties, unconventional magnets have recently garnered attention in the context of superconducting spintronics by inducting high-parities spin-triplet states [34, 45, 46, 47, 48, 49], orientation-dependent transport in altermagnetic-supercondctor hybridstructure [50, 51, 21, 52, 53, 54, 20, 21, 55], gate-controlled Josephson junctions [46, 56, 57, 58, 59, 60], topological superconductors [61, 49, 62, 63, 64, 65], nonreciprocal field-free superconducting devices [66, 67, 68] and ther superconducting phenomena generated by the interplay between unconventional magnetism and superconductivity [11]. Together, these developments establish unconventional magnets as a fertile ground for exploring spintronic functionalities without net magnetization.
@@
In this work, we explore spin-selective transport in heterostructures composed of AMs and normal metals, aiming to establish all-electrical spintronic functionality that harnesses the intrinsic, nonrelativistic spin splitting of AMs. Two AM-based spintronics devices with fundamental interest are proposed, including a spin filter and spin valve (Fig. 1). In the regime of quantum-coherent transport, a conventional spin filter is realized in a ferromagnet junction, where the Zeeman splitting in combination with a build-up potential barrier gives rise to a spin-polarized flow since spin-up and spin-down electrons experience different barrier heights [73, 12, 13]. While the spin polarization of the current through the junction is determined by the magnetization of the ferromagnet, the applied barrier can tune the allowed occupation of states belonging to the lower Zeeman level and thus strengthen the spin filtering. In the proposed spin filter [Fig. 1(a)], a gate applied to the AM region generates switchable fully spin-polarized conductance by selectively blocking one spin channel, yet without requiring a Zeeman exchange field. Furthermore, two AM-based spin filters connected in series compose a spin valve [Fig. 1(b)]. This is similar to the conventional spin valve composed of a conventional magnetic bilayer, where only electrons with spin aligned to the magnetization contribute to the current [13, 12]. Consequently, conductance is high when the magnetization directions of the two ferromagnetic layers are aligned, while it is suppressed when these directions are opposite, known as the parallel and antiparallel configurations, respectively. Parallel-antiparallel switching is typically achieved by applying an external magnetic field, which controls the relative magnetization orientations between layers. The resulting difference in conductance and thus the resistance under field reversal gives rise to the giant magnetoresistance effect [74, 75], which forms the basis for modern electronics applications such as data storage technologies, magnetic random access memory, and magnetic field sensors [12, 13, 14, 15, 16]. The behaviors analogous to the parallel-antiparallel switching are also found in the proposed spin valve based on strong AMs, where the alteration of the spin polarization in each spin filter is achieved by the applied gate [Fig. 1(b)], without an external magnetic field and net magnetization.
@@
The remainder of this work is organized as follows. The band structure, Fermi surface features of weak and strong AM are demonstrated in Sec. II, which determines the transport properties and the spin filter effect in the altermagnetic junction illustrated in Sec. III. Based on these properties, two spintronic devices, gate-controlled spin filter and spin valve, are proposed in Sec. IV. The conclusion is given in Sec. V.
II Weak and strong altermagnets
We start by introducing the Hamiltonian of $d$-wave AMs in the spin basis $\left(\psi_{\bm{k},\uparrow},\psi_{\bm{k},\downarrow}\right)^{\mathrm{T}}$ [7, 8]
@@ II Weak and strong altermagnets
for a fixed energy $E$. These open Fermi surfaces impose directional constraints on the availability of spin-polarized electronic states. As shown in the inset of Fig. 2(b), for a given positive energy $E$, spin-up (spin-down) states are absent when $a|k_{x}|<\sqrt{E/(t+J)}$ [$a|k_{y}|<\sqrt{E/(t+J)}$], due to the absence of real solutions for the corresponding hyperbolic Fermi contour. Conversely, for negative energies $E<0$, spin-up (spin-down) states are absent when $a|k_{y}|<\sqrt{|E|/(t-J)}$ [$a|k_{x}|<\sqrt{|E|/(t-J)}$]. This directional open Fermi surface is a hallmark of the strong altermagnetic regime and stands in sharp contrast to the weak AM case, where both spin species coexist over closed elliptical Fermi surfaces. While our discussions above are based on $d_{x^{2}-y^{2}}$-wave AM with $\theta_{J}=0$, the distinct properties such as the dispersions, Fermi surface geometry between weak and strong AM, are shared in all $d$-wave AM with arbitrary $\theta_{J}$.
III Spin-Selective Transport in Altermagnetic Junctions

### v3

@@ Abstract
The recent discovery of altermagnets, which exhibit spin splitting without net magnetization, opens new directions for spintronics beyond the limits of ferromagnets, antiferromagnets, and spin–orbit coupled systems. We investigate spin-selective quantum transport in heterostructures composed of a normal metal and a two-dimensional $d$-wave altermagnet, and identify a universal mechanism for achieving perfect spin polarization. The mechanism is dictated by Fermi-surface geometry: closed surfaces in weak altermagnets yield partial and oscillatory spin filtering, whereas open surfaces in strong altermagnets intrinsically enforce fully spin-polarized conductance. Exploiting these distinct transport properties, we propose all-electrical spin-filter and spin-valve architectures, where resonant tunneling produces highly spin-polarized conductance tunable by gate voltage and interface transparency. Altermagnets with open Fermi surfaces further support gate-reversible perfect spin polarization that remains robust against interface scattering, disorder, and temperature. We also demonstrate an electrically controlled spin valve that reproduces the functionality of magnetic tunnel junctions without magnetic fields or relativistic mechanisms. $d$-wave altermagnets with open Fermi surfaces thus provide a new platform for low-dissipation, scalable, and magnetic-field-free spintronic devices, with potential for integration into next-generation quantum and CMOS-compatible technologies.
Introduction
@@
Unconventional magnets have emerged as a new class of systems that lie beyond the traditional dichotomy of ferromagnets and antiferromagnets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. These so-called third type of magnets support spin-split Fermi surfaces resembling those of ferromagnets [12, 13], yet exhibiting globally vanishing magnetization due to a compensated magnetic ordering, akin to antiferromagnets [14, 15, 16]. Depending on the momentum-space symmetry of the exchange field and the underlying crystalline symmetries, unconventional magnets can be classified according to the angular character of the spin splitting, including $p$-, $d$-, $f$-, $g$-, and $i$-wave types [7, 8, 9, 17, 18, 19, 20, 21, 22, 23]. Among these, $d$-, $g$-, and $i$-wave magnets are known as altermagnets (AMs), characterized by simultaneously broken time-reversal symmetry and spatial rotational symmetry, but preserving their joint symmetry [7, 8]. Altermagnets exhibit fully compensated spins and broken parity-time ($\bm{PT}$) symmetry. [3]. Furthermore, AMs can also be classified according to their closed and open Fermi surfaces, [7, 8, 24, 25, 26], leading to the concept of weak and strong AMs. In contrast, $p$- and $f$-wave magnets break only rotational symmetry but preserve time-reversal symmetry originating from nonsymmorphic lattice symmetries [9, 10], typically exhibiting closed anisotropic Fermi surfaces.
@@
Unconventional magnets are promising platforms for spintronic devices [13, 14]. The anisotropic spin split Fermi surface offers mechanisms for giant magnetoresistance [27], spin-orbit-free anomalous Hall effects [28, 1, 29, 30, 31], spin-transfer torques [32, 33, 34], spin filtering effects [35], spin pumping effects [36], non-linear transports [21], light-matter interactions [37, 38], light-induced spin density [39], non-Hermitian electronic responses [40, 41], strongly correlation in Mott insulators [42] and other novel phenomena [43, 44]. These mechanisms are expected in experimentally accessible altermagnetic materials exhibiting novel spin transport, such as RuO2 [45, 28, 30], MnTe [46, 29, 31, 47], Mn5Si3 [31, 47], CrSb [48, 49], FeS2 [50], MnF2 [51], Mn2Au [52], La2O3Mn2Se2, Ba2CaOsO6 [53] and La2CuO4 [54].
In addition to their normal-state properties, unconventional magnets have recently garnered attention in the context of superconducting spintronics by inducting high-parities spin-triplet states [39, 55, 56, 57, 58, 59], orientation-dependent transport in altermagnetic-supercondctor hybridstructure [60, 61, 26, 62, 63, 64, 25, 26, 65], gate-controlled Josephson junctions [56, 66, 67, 68, 69, 70], topological superconductors [71, 59, 72, 73, 74, 75], nonreciprocal field-free superconducting devices [76, 77, 78] and their superconducting phenomena generated by the interplay between unconventional magnetism and superconductivity [11]. Together, these developments establish unconventional magnets as a fertile ground for exploring spintronic functionalities without net magnetization.
@@
In this work, we explore spin-selective transport in heterostructures composed of AMs and normal metals, aiming to establish all-electrical spintronic functionality that harnesses the intrinsic, non-relativistic spin splitting of AMs. Two AM-based spintronics devices with fundamental interest are proposed, including a spin filter and spin valve (Fig. 1). In the regime of quantum-coherent transport, a conventional spin filter is realized in a ferromagnet junction, where the Zeeman splitting in combination with a build-up potential barrier gives rise to a spin-polarized flow since spin-up and spin-down electrons experience different barrier heights [83, 12, 13]. While the spin polarization of the current through the junction is determined by the magnetization of the ferromagnet, the applied barrier can tune the allowed occupation of states belonging to the lower Zeeman level and thus strengthen the spin filtering. In the proposed spin filter [Fig. 1(a)], a gate applied to the AM region generates switchable fully spin-polarized conductance by selectively blocking one spin channel, yet without requiring a Zeeman exchange field. Furthermore, two AM-based spin filters connected in series compose a spin valve [Fig. 1(b)]. This is similar to the conventional spin valve composed of a conventional magnetic bilayer, where only electrons with spin aligned to the magnetization contribute to the current [13, 12]. Consequently, conductance is high when the magnetization directions of the two ferromagnetic layers are aligned, while it is suppressed when these directions are opposite, known as the parallel and antiparallel configurations, respectively. Parallel-antiparallel switching is typically achieved by applying an external magnetic field, which controls the relative magnetization orientations between layers. The resulting difference in conductance and thus the resistance under field reversal gives rise to the giant magnetoresistance effect [84, 85], which forms the basis for modern electronics applications such as data storage technologies, magnetic random access memory, and magnetic field sensors [12, 13, 14, 15, 16]. The behaviors analogous to the parallel-antiparallel switching are also found in the proposed spin valve based on strong AMs, where the alteration of the spin polarization in each spin filter is achieved by the applied gate [Fig. 1(b)], without an external magnetic field and net magnetization.
@@
Results
Weak and strong altermagnets
We start by introducing the Hamiltonian of $d$-wave AMs in continuum limit in the spin basis $\left(\psi_{\bm{k},\uparrow},\psi_{\bm{k},\downarrow}\right)^{\mathrm{T}}$ [7, 8]
@@ Results
for a fixed energy $E$. These open Fermi surfaces impose directional constraints on the availability of spin-polarized electronic states. As shown in the inset of Fig. 2(b), for a given positive energy $E$, spin-up (spin-down) states are absent when $a|k_{x}|<\sqrt{E/(t+J)}$ $a|k_{y}|<\sqrt{E/(t+J)}$], due to the absence of real solutions for the corresponding hyperbolic Fermi contour. Conversely, for negative energies $E<0$, spin-up (spin-down) states are absent when $a|k_{y}|<\sqrt{|E|/(t-J)}$ $a|k_{x}|<\sqrt{|E|/(t-J)}$]. This directional open Fermi surface is a hallmark of the strong altermagnetic regime and stands in sharp contrast to the weak AM case, where both spin species coexist over closed elliptical Fermi surfaces. While our discussions above are based on $d_{x^{2}-y^{2}}$-wave AM with $\theta_{J}=0$, the distinct properties such as the dispersions, Fermi surface geometry between weak and strong AM, are shared in all $d$-wave AM with arbitrary $\theta_{J}$. Remarkably, the distinct Fermi-surface geometries described by the continuum altermagnetic Hamiltonian [Eq. 1] also emerge in lattice-model Hamiltonian fitted to real materials such as RuO2 [45, 28, 30], La2O3Mn2Se2, and Ba2CaOsO6 [53] (see Sec. I of the Supplemental Material). In the following, we therefore restrict our investigation to the continuum model, which clearly illustrates the underlying mechanism of spin-selective transport. This mechanism is further corroborated by the lattice model results presented in the Discussion section.
Spin-Selective Transport in Altermagnetic Junctions